# Generative quantum machine learning via denoising diffusion probabilistic models

### 01 Feb 2024 v3

@@ Abstract
Deep generative models are key-enabling technology to computer vision, text generation and large language models. Denoising diffusion probabilistic models (DDPMs) have recently gained much attention due to their ability to generate diverse and high-quality samples in many computer vision tasks, as well as to incorporate flexible model architectures and relatively simple training scheme. Quantum generative models, empowered by entanglement and superposition, have brought new insight to learning classical and quantum data. Inspired by the classical counterpart, we propose the quantum denoising diffusion probabilistic model (QuDDPM) to enable efficiently trainable generative learning of quantum data. QuDDPM adopts sufficient layers of circuits to guarantee expressivity, while introduces multiple intermediate training tasks as interpolation between the target distribution and noise to avoid barren plateau and guarantee efficient training. We provide bounds on the learning error and demonstrate QuDDPM’s capability in learning correlated quantum noise model, quantum many-body phases and topological structure of quantum data. The results provide a paradigm for versatile and efficient quantum generative learning.
@@
Variational parameterized quantum circuits (PQC) [1, 2, 3, 4] provide a near-term platform for quantum machine learning [5, 6, 7]. In terms of generative models [8, 9, 10, 11], quantum generative adversarial networks (QuGANs) is recently proposed [12, 13, 14, 15, 16], in analogy to classical generative adversarial networks (GANs) [17]. Despite the success, classical GAN models are known for training issues such as mode collapse. In classical deep learning, denoising diffusion probabilistic models (DDPMs) and their close relatives [18, 19, 20, 21, 22] have recently gained much attention due to relatively simple training schemes and their ability to generate diverse and high-quality samples in many computer vision tasks [23, 24, 25, 26] over the best GANs, and to incorporate flexible model architectures.
@@
Figure 1: Schematic of QuDDPM. The forward noisy process is implemented by quantum scrambling circuit (QSC) in (a), while in the backward denoising process is achieved via measurement enabled by ancilla and PQC in (d). Subplots (b1-b5) and (c1-c4) presents the Bloch sphere dynamics in generation of states clustering around |$\ket{0}$, where convergence can be seen despite sample fluctuations (see Appendix K).
@@
In this work, we propose the quantum denoising diffusion probabilistic model (QuDDPM) as an efficiently trainable scheme to generative quantum learning, through a coordination between a forward noisy diffusion process via quantum scrambling [27, 28] and a backward denoising process via quantum measurement. We provide bounds on the learning error and then demonstrate QuDDPM’s capability in examples relevant to characterizing quantum device noises, learning quantum many-body phases and capturing topological structure of quantum data. For an $n$-qubit problem, QuDDPM adopts linear-in-$n$ layers of circuits to guarantee expressivity, while introduces $T\sim n/\log(n)$ intermediate training tasks and to guarantee efficient training.
@@ I General formulation of QuDDPM
We consider the task of the generating new elements from an unknown distribution ${\cal E}_{0}$ of quantum states, provided only a number of samples ${\cal S}_{0}=\{\ket{\psi_{k}}\}\sim{\cal E}_{0}$ from the distribution. The task under consideration—generating individual states from the distribution (e.g., a single Haar random state or $K$-design state)—is not equivalent to generating the average state of a distribution (e.g., a fully mixed state for Haar ensemble) considered in previous works of QuGAN [12]. To complete the task, QuDDPM learns a map from a noisy unstructured distribution of states to the structured target distribution ${\cal E}_{0}$. It does so via a divide-and-conquer strategy of creating smooth interpolations between the target distribution and full noise, so that the training is divided to sub-tasks on low-depth circuit to avoid barren plateau [29, 30, 31, 32].
@@
As shown in Fig. 1, QuDDPM includes two quantum circuits, one to enable the forward diffusion of sample data towards noise via scrambling and one to realize the backward denoising from noise towards generated data via measurement. For each data |$\ket{\psi_{i}^{(0)}}$, the forward scrambling circuit (Fig. 1a) samples a series of $T$ random unitary gates $U_{1}^{(i)},\dots,U^{(i)}_{T}$ independently, such that the ensemble ${\cal S}_{k}=\{\ket{\psi_{i}^{(k)}}=\prod_{\ell=1}^{k}U_{\ell}^{(i)}\ket{\psi_{i}^{(0)}}\}_{i}$ evolves from the sample data towards a random ensemble of pure states from $k=0$ to $k=T$. A Bloch sphere visualization of such a forward scrambling dynamics is depicted in Fig. 1(b1-b5) for a toy problem of learning single-qubit states ${\cal S}_{0}$ clustered around a single pure state, e.g. |$\ket{0}$, (b1), where the noise ${\cal S}_{T}$ is uniform on Bloch sphere (b5).
@@
With the interpolation from data ${\cal S}_{0}$ and noise ${\cal S}_{T}$ in hand, the backward process can start from randomly sampled noise $\tilde{S}_{T}$ (Fig. 1c5) and reduce the noise gradually via measurement step by step, towards the final generated data $\tilde{{\cal S}}_{0}$ (Fig. 1c1) that mimic the sample data (Fig. 1b1). Measurements are necessary, as the denoising map is contractive and maintain the purity of each generated data in $\tilde{{\cal S}}_{0}$. As shown in Fig. 1d, each denoising step adopts a unitary $\tilde{U}_{k}$ on $n_{A}$ number of ancilla qubits in |$\ket{0}$ and performs a projective measurement in computational bases on the ancilla after the unitary $\tilde{U}_{k}$. Starting from the state |$\ket{\tilde{\psi}_{i}^{(T)}}$, which is randomly sampled from noise ensemble, each unitary plus measurement step evolves the random state towards the generated data |$\ket{\tilde{\psi}_{i}^{(0)}}$. Note that here all unitaries $\tilde{U}_{k}$ are fixed after training. In practice, the generation of noisy |$\ket{\tilde{\psi}_{i}^{(T)}}$ can be directly completed by running the $T$ layers of forward scrambling circuit. Via training, the denoising process learns information about target from the ensembles in the forward scrambling, stores information in the circuit parameters, and then encodes onto the generated data.
@@ II Training strategy
In classical DDPM, the Gaussian nature of the diffusion allows efficient training via maximizing an evidence lower bound (ELBO) for the log-likelihood function, which can be evaluated analytically [18, 19] (see Appendix E). However, in QuDDPM, we do not expect such analytical simplification to exist at all—classical simulation of quantum device is inherently inefficient. Instead, the training of the QuDDPM relies on the capability of quantum measurements to extract information about ensemble of quantum states for the efficient evaluation of a loss function.
The training of a $T$-step QuDDPM consists of $T$ training cycles, starting from the first denoising step $\tilde{U}_{T}$ towards the last $\tilde{U}_{1}$. As shown in Fig. 2, at the training cycle $(T+1-k)$, the forward noisy diffusion process is implemented from $U_{1}^{(i)}$ to $U_{k}^{(i)}$ to generates the noisy ensemble ${\cal S}_{k}=\{\ket{\psi_{i}^{(k)}}\}_{i}$ While the backward denoising process performs the denoising steps $\tilde{U}_{T}$ to $\tilde{U}_{k+1}$ to generate the denoising ensemble $\tilde{{\cal S}}_{k}=\{\ket{\tilde{\psi}_{i}^{(k)}}\}_{i}$. Within the training cycle, the parameters of the denoising PQC $\tilde{U}_{k+1}$ are updated such that the generated denoising ensemble $\tilde{{\cal S}}_{k}$ converges to the noisy ensemble ${\cal S}_{k}$. Therefore, QuDDPM divides the original training problem into $T$ smaller and easier ones. Indeed, even with a global loss function, we can divide the $\Omega(n)$ layers (required by expressivity) of gates to $T\in\Omega(n/\log n)$ diffusion steps, such that each $\tilde{U}_{k+1}$ has order $\log(n)$ layers of gates to avoid barren plateau [30].
@@ III Loss function
To enable training, a loss function quantifies the distance between the two ensembles of quantum states. In this work, we focus on the Maximum Mean Discrepancy (MMD) [33] (see Appendix F) and the Wasserstein distance [34, 35] (see Appendix B) based on the state overlaps $\lvert\braket{\psi_{i}^{(k)}}{\tilde{\psi}_{i}^{(k)}}\rvert^{2}$ estimated via a swap test (see Appendix G).
Given two independent distributions of pure states ${\cal E}_{1}$ and ${\cal E}_{2}$ on the state vector space $V$. The state-wise fidelity between |$\ket{\psi}$ and |$\ket{\phi}$ is defined as $F(\ket{\phi},\ket{\psi})=\lvert\braket{\phi}{\psi}\rvert^{2}$, and we can further define the mean fidelity as
@@
where the random states |$\ket{\phi}\sim{\cal E}_{1}$ and |$\ket{\psi}\sim{\cal E}_{2}$ are drawn independently. Since the fidelity $F$ is a symmetric and positive definite quadratic kernel, according to the theory of reproducing kernel Hilbert space (RKHS), the MMD distance can be written as (see Appendix F)
@@
which allows the estimation of MMD through sampled state ensembles ${\cal S}_{1}$ and ${\cal S}_{2}$. The expressivity of general MMD as a statistical distance measure depends on the kernel. On one hand, identifiability requires, distance is zero if and only if ${\cal E}_{1}={\cal E}_{2}$. On the other hand, one also needs to ensure the quality of statistical estimation of the distance with finite sample size of state ensembles. Hence, whether fidelity (1) is a proper kernel choice is problem-dependent. In Appendix I, we show an example where ${\cal D}_{\text{MMD}}$ in Eq. (2) fails to distinguish two simple distributions. To resolve such an issue, we may alternatively consider the Wasserstein distance, a geometrically meaningful distance for comparing complex data distributions based on the theory of optimal transport [34, 35] (see Appendix B for details).
Figure 3: The decay of MMD distance ${\cal D}$ between generated ensemble $\tilde{{\cal S}}_{t}$ using different models and target ensemble of states ${\cal E}_{0}$ clustered around |$\ket{0,0}$ versus training steps. The converged value is ${\cal D}\simeq 0.002$ for QuDDPM, showing a two-order-of-magnitude advantage over QuDT and QuGAN.
@@
where ${\cal D}$ can be the MMD distance or the Wasserstein distance. The distribution ~$\tilde{{\cal E}}_{k}$ is a function of all the reverse denoising steps and the noise distribution ~$\tilde{{\cal E}}_{T}$. In practice, we use finite samples to approximate the loss function as (${\cal L}(\{\tilde{U}_{\ell}\}_{k+1}^{T},{\cal S}_{k},\tilde{{\cal S}}_{T})$.
The toy example in Fig. 1 adopted the MMD distance in the loss function and details of the training can be found in Appendix K (see Ref. [36] for codes and data). Here we present the training history of a more challenging two-qubit example of preparing states clustered around |$\ket{0,0}$, to allow a meaningful comparison with other algorithms. In each of the 20 steps of training cycles, the loss function is minimized till convergence. To quantify the convergence, we also evaluate the MMD distance (see Fig. 3) ${\cal D}(\tilde{{\cal S}}_{t},{\cal E}_{0})$ between the true distribution ${\cal E}_{0}$ and the trained ensemble of states $\tilde{{\cal S}}_{t}$ throughout the training cycles (blue), showing a convergence towards ${\cal D}=0$. The periodic spikes shows the initial increase of the MMD distance at each training cycle, due to introducing randomly initialized PQC in a new denoising step. For reference, we also plot the evolution of the MMD distance throughout the forward-diffusion (red circles), which starts from zero at diffusion step 0 and grows towards larger value as diffusion step increases (from right to left). We see the training results (blue) follows closely to the diffusion results (red) as expected. In addition, the testing results (green) also agree well with the training results (blue) for QuDDPM.
As benchmarks, we consider two major quantum generative models, QuGAN and quantum direct transport (QuDT). QuDT can be regarded as the generalization of quantum circuit Born machine [37, 38, 39, 40] towards quantum data. Previous works of both models only considered a single quantum state or classical distributions [14, 41, 38], and here we generalize them to adapt to the state ensemble generation task by allowing Haar random states as inputs and introducing ancilla to be measured (see Appendix L). For a fair comparison, we keep the number of variational parameters of generator circuits in QuDT and QuGAN the same as QuDDPM, listed in Tab. 2. As shown in Fig. 3, QuDT and QuGAN converges to generate ensembles with a substantial MMD deviation to the true ensemble, demonstrating QuDDPM’s advantage due to its unique diffusion and denoising process.
@@ IV Gate complexity and convergence
Now we discuss about the number of local gates required and convergence analysis for QuDDPM to solve an $n$-qubit generative task. For simplicity, we assume the qubits are one-dimensional with nearest neighbor interactions, while similar counting can be done for other cases. To guarantee convergence towards noise, the forward scrambling circuits needs number of layers linear in $n$ as predicted by $K$-design [42, 43], leading to $O(n^{2})$ total gates. The backward circuit will be similar, with at most $n_{A}\leq 2n$ additional ancillas and $O(n^{2})$ gates, leading to the overall gate complexity of QuDDPM to be $O(n^{2})$.
@@
with deviation of ${\cal S}_{T}$ to true random states $E_{\rm diff}$, measurement error $E_{{\cal M}}$ and generalization error $E_{\rm gen}$. We discuss the scaling of three part separately in the following. Suppose the diffusion circuits approach approximate $K$-design, its diffusion error is known as [42]
@@
has the scaling $\mathcal{O}\left(\frac{1}{TN}\right)$, as shown in Fig. 4 for a $n=4$ qubit clustering state generation task. Here we estimate the generalization error via a validation set independently sampled, while the proof is an open problem [45, 46]. The $1/N$ scaling agrees with classical machine learning results [47, 48].

### 16 Feb 2024 v4

@@ Abstract
Deep generative models are key-enabling technology to computer vision, text generation, and large language models. Denoising diffusion probabilistic models (DDPMs) have recently gained much attention due to their ability to generate diverse and high-quality samples in many computer vision tasks, as well as to incorporate flexible model architectures and a relatively simple training scheme. Quantum generative models, empowered by entanglement and superposition, have brought new insight to learning classical and quantum data. Inspired by the classical counterpart, we propose the quantum denoising diffusion probabilistic model (QuDDPM) to enable efficiently trainable generative learning of quantum data. QuDDPM adopts sufficient layers of circuits to guarantee expressivity, while it introduces multiple intermediate training tasks as interpolation between the target distribution and noise to avoid barren plateau and guarantee efficient training. We provide bounds on the learning error and demonstrate QuDDPM’s capability in learning correlated quantum noise model, quantum many-body phases, and topological structure of quantum data. The results provide a paradigm for versatile and efficient quantum generative learning.
@@
Variational parametrized quantum circuits (PQCs) [1, 2, 3, 4] provide a near-term platform for quantum machine learning [5, 6, 7]. In terms of generative models [8, 9, 10, 11], quantum generative adversarial networks (QuGANs) have been recently proposed [12, 13, 14, 15, 16], in analogy to classical generative adversarial networks (GANs) [17]. Despite the success, classical GAN models are known for training issues such as mode collapse. In classical deep learning, denoising diffusion probabilistic models (DDPMs) and their close relatives [18, 19, 20, 21, 22] have recently gained much attention due to relatively simple training schemes and their ability to generate diverse and high-quality samples in many computer vision tasks [23, 24, 25, 26] over the best GANs, and to incorporate flexible model architectures.
@@
Figure 1: Schematic of QuDDPM. The forward noisy process is implemented by a quantum scrambling circuit (QSC) in (a), while in the backward denoising process is achieved via measurement enabled by ancilla and PQC in (d). Subplots (b1)-(b5) and (c1)-(c5) present the Bloch sphere dynamics in generation of states clustering around |$\ket{0}$, where convergence can be seen despite sample fluctuations, as shown in the Appendix L.
@@
In this work, we propose the quantum denoising diffusion probabilistic model (QuDDPM) as an efficiently trainable scheme to generative quantum learning, through a coordination between a forward noisy diffusion process via quantum scrambling [27, 28] and a backward denoising process via quantum measurement. We provide bounds on the learning error and then demonstrate QuDDPM’s capability in examples relevant to characterizing quantum device noises, learning quantum many-body phases and capturing topological structure of quantum data. For an $n$-qubit problem, QuDDPM adopts linear-in-$n$ layers of circuits to guarantee expressivity, while it introduces $T\sim n/\log(n)$ intermediate training tasks to guarantee efficient training.
@@ I General formulation of QuDDPM
We consider the task of the generating new elements from an unknown distribution ${\cal E}_{0}$ of quantum states, provided only a number of samples ${\cal S}_{0}=\{\ket{\psi_{k}}\}\sim{\cal E}_{0}$ from the distribution. The task under consideration—generating individual states from the distribution (e.g., a single Haar random state or $K$-design state)—is not equivalent to generating the average state of a distribution (e.g., a fully mixed state for Haar ensemble) considered in previous works of QuGAN [12]. To complete the task, QuDDPM learns a map from a noisy unstructured distribution of states to the structured target distribution ${\cal E}_{0}$. It does so via a divide-and-conquer strategy of creating smooth interpolations between the target distribution and full noise, so that the training is divided to subtasks on a low-depth circuit to avoid barren plateau [29, 30, 31, 32].
@@
As shown in Fig. 1, QuDDPM includes two quantum circuits, one to enable the forward diffusion of sample data toward noise via scrambling and one to realize the backward denoising from noise toward generated data via measurement. For each data |$\ket{\psi_{i}^{(0)}}$, the forward scrambling circuit [Fig. 1 (a)] samples a series of $T$ random unitary gates $U_{1}^{(i)},\dots,U^{(i)}_{T}$ independently, such that the ensemble ${\cal S}_{k}=\{\ket{\psi_{i}^{(k)}}=\prod_{\ell=1}^{k}U_{\ell}^{(i)}\ket{\psi_{i}^{(0)}}\}_{i}$ evolves from the sample data toward a random ensemble of pure states from $k=0$ to $k=T$. A Bloch sphere visualization of such a forward scrambling dynamics is depicted in Figs. 1(b1)-(b5) for a toy problem of learning single-qubit states ${\cal S}_{0}$ clustered around a single pure state, e.g. |$\ket{0}$, (b1), where the noise ${\cal S}_{T}$ is uniform on the Bloch sphere (b5).
@@
With the interpolation from data ${\cal S}_{0}$ and noise ${\cal S}_{T}$ in hand, the backward process can start from randomly sampled noise $\tilde{S}_{T}$ [Fig. 1(c5)] and reduce the noise gradually via measurement step by step, toward the final generated data $\tilde{{\cal S}}_{0}$ [Fig. 1(c1)] that mimic the sample data [Fig. 1(b1)]. Measurements are necessary, as the denoising map is contractive and maintains the purity of all generated data in $\tilde{{\cal S}}_{0}$. As shown in Fig. 1(d), each denoising step adopts a unitary $\tilde{U}_{k}$ on the system plus $n_{A}$ number of ancilla qubits in |$\ket{0}$ and performs a projective measurement in computational bases on the ancilla after the unitary $\tilde{U}_{k}$. Starting from the state |$\ket{\tilde{\psi}_{i}^{(T)}}$, which is randomly sampled from noise ensemble, each unitary plus measurement step evolves the random state toward the generated data |$\ket{\tilde{\psi}_{i}^{(0)}}$. Note that here all unitaries $\tilde{U}_{k}$ are fixed after training. In practice, the generation of noisy |$\ket{\tilde{\psi}_{i}^{(T)}}$ can be directly completed by running the $T$ layers of the forward scrambling circuit on an arbitrary initial state. Via training, the denoising process learns information about the target from the ensembles in the forward scrambling, stores information in the circuit parameters, and then encodes onto the generated data.
@@ II Training strategy
In classical DDPM, the Gaussian nature of the diffusion allows efficient training via maximizing an evidence lower bound for the log-likelihood function, which can be evaluated analytically [18, 19] (see Appendix E). However, in QuDDPM, we do not expect such analytical simplification to exist at all—classical simulation of quantum device is inherently inefficient. Instead, the training of the QuDDPM relies on the capability of quantum measurements to extract information about the ensemble of quantum states for the efficient evaluation of a loss function.
The training of a $T$-step QuDDPM consists of $T$ training cycles, starting from the first denoising step $\tilde{U}_{T}$ toward the last $\tilde{U}_{1}$. As shown in Fig. 2, at the training cycle $(T+1-k)$, the forward noisy diffusion process is implemented from $U_{1}^{(i)}$ to $U_{k}^{(i)}$ to generate the noisy ensemble ${\cal S}_{k}=\{\ket{\psi_{i}^{(k)}}\}_{i}$, While the backward denoising process performs the denoising steps $\tilde{U}_{T}$ to $\tilde{U}_{k+1}$ to generate the denoising ensemble $\tilde{{\cal S}}_{k}=\{\ket{\tilde{\psi}_{i}^{(k)}}\}_{i}$. Within the training cycle, the parameters of the denoising PQC $\tilde{U}_{k+1}$ are updated such that the generated denoising ensemble $\tilde{{\cal S}}_{k}$ converges to the noisy ensemble ${\cal S}_{k}$. Therefore, QuDDPM divides the original training problem into $T$ smaller and easier ones. Indeed, even with a global loss function for $n$ qubits, we can divide the $\Omega(n)$ layers (required by expressivity) of gates into $T\in\Omega(n/\log n)$ diffusion steps, such that each $\tilde{U}_{k+1}$ has order $\log(n)$ layers of gates to avoid barren plateau [30].
@@ III Loss function
To enable training, a loss function quantifies the distance between the two ensembles of quantum states. In this work, we focus on the maximum mean discrepancy (MMD) [45] (see Appendix F) and the Wasserstein distance [46, 47] (see Appendix B) based on the state overlaps $\lvert\braket{\psi_{i}^{(k)}}{\tilde{\psi}_{i}^{(k)}}\rvert^{2}$ estimated via a swap test (see Appendix H).
Given two independent distributions of pure states ${\cal E}_{1}$ and ${\cal E}_{2}$ on the state vector space $V$, the statewise fidelity between |$\ket{\psi}$ and |$\ket{\phi}$ is defined as $F(\ket{\phi},\ket{\psi})=\lvert\braket{\phi}{\psi}\rvert^{2}$, and we can further define the mean fidelity as
@@
where the random states |$\ket{\phi}\sim{\cal E}_{1}$ and |$\ket{\psi}\sim{\cal E}_{2}$ are drawn independently. Since the fidelity $F$ is a symmetric and positive definite quadratic kernel, according to the theory of reproducing kernel Hilbert space [61], the MMD distance can be written as (see Appendix F)
@@
which allows the estimation of MMD through sampled state ensembles ${\cal S}_{1}$ and ${\cal S}_{2}$. The expressivity of general MMD as a statistical distance measure depends on the kernel. On one hand, identifiability requires that the distance be zero if and only if ${\cal E}_{1}={\cal E}_{2}$. On the other hand, one also needs to ensure the quality of statistical estimation of the distance with a finite sample size of state ensembles. Hence, whether fidelity (1) is a proper kernel choice is problem-dependent. In Appendix J, we show an example where ${\cal D}_{\text{MMD}}$ in Eq. (2) fails to distinguish two simple distributions. To resolve such an issue, we may alternatively consider the Wasserstein distance, a geometrically meaningful distance for comparing complex data distributions based on the theory of optimal transport [46, 47] (see the Appendix B for details).
Figure 3: The decay of MMD distance ${\cal D}$ between generated ensemble $\tilde{{\cal S}}_{t}$ using different models and target ensemble of states ${\cal E}_{0}$ clustered around |$\ket{0,0}$ versus training steps. The converged value is ${\cal D}\simeq 0.002$ for QuDDPM, showing an advantage of $2$ orders of magnitude over QuDT and QuGAN.
@@
where ${\cal D}$ can be the MMD distance or the Wasserstein distance. The distribution ~$\tilde{{\cal E}}_{k}$ is a function of all the reverse denoising steps from $T$ to $k+1$ and the noise distribution ~$\tilde{{\cal E}}_{T}$. In practice, we use finite samples to approximate the loss function as (${\cal L}(\{\tilde{U}_{\ell}\}_{k+1}^{T},{\cal S}_{k},\tilde{{\cal S}}_{T})$.
The toy example in Fig. 1 adopted the MMD distance in the loss function and details of the training can be found in Appendix L (see Ref. [48] for codes and data, and Table 1 for details of parameters). Here we present the training history of a more challenging $2$-qubit example of preparing states clustered around |$\ket{0,0}$, to allow a meaningful comparison with other algorithms. In each of the 20 steps of training cycles, the loss function is minimized till convergence. To quantify the convergence, we also evaluate the MMD distance (see Fig. 3) ${\cal D}(\tilde{{\cal S}}_{t},{\cal E}_{0})$ between the true distribution ${\cal E}_{0}$ and the trained ensemble of states $\tilde{{\cal S}}_{t}$ throughout the training cycles (blue), showing a convergence toward ${\cal D}=0$. The periodic spikes show the initial increase of the MMD distance at each training cycle, due to introducing a randomly initialized PQC in a new denoising step. For reference, we also plot the evolution of the MMD distance throughout the forward-diffusion (red circles), which starts from zero at diffusion step 0 and grows toward a larger value as the diffusion step increases (from right to left). We see the training results (blue) follow closely to the diffusion results (red) as expected. In addition, the testing results (green) also agree well with the training results (blue) for QuDDPM.
As benchmarks, we consider two major quantum generative models, QuGAN and quantum direct transport (QuDT). QuDT can be regarded as the generalization of the quantum circuit Born machine [49, 50, 51, 52] toward quantum data. Previous works of both models only considered a single quantum state or classical distributions [14, 53, 50], and here we generalize them to adapt to the state ensemble generation task by allowing Haar random states as inputs and introducing ancilla to be measured (see Appendix M). For a fair comparison, we keep the number of variational parameters of generator circuits in QuDT and QuGAN the same as QuDDPM, listed in Tabble 2. As shown in Fig. 3, QuDT and QuGAN converge to generate ensembles with a substantial MMD deviation to the true ensemble, demonstrating QuDDPM’s advantage due to its unique diffusion and denoising process.
@@ IV Gate complexity and convergence
Now we discuss the number of local gates required and convergence analysis for QuDDPM to solve an $n$-qubit generative task. For simplicity, we assume the qubits are one-dimensional with nearest-neighbor interactions, while similar counting can be done for other cases. To guarantee convergence toward noise, the forward scrambling circuits need a linear number of layers in $n$ as predicted by $K$ design [54, 55], leading to $O(n^{2})$ total gates. The backward circuit will be similar, with at most $n_{A}\leq 2n$ additional ancillas and $O(n^{2})$ gates, leading the overall gate complexity of QuDDPM to be $O(n^{2})$.
@@
with a deviation $E_{\rm diff}$ of ${\cal S}_{T}$ to true random states, measurement error $E_{{\cal M}}$ and generalization error $E_{\rm gen}$. We discuss the scaling of three parts separately in the following. Suppose the diffusion circuits approach an approximate $K$ design; its diffusion error is known as [54]
@@
has the scaling $\mathcal{O}\left(\frac{1}{TN}\right)$, as shown in Fig. 4 for an $n=4$ qubit clustering state generation task. Here we estimate the generalization error via a validation set independently sampled, while the proof is an open problem [57, 58]. The $1/N$ scaling agrees with classical machine learning results [59, 60].